The price of a European call option on a non-dividend-paying stock with a strike price of $50 is $6. The stock price is $51, the continuously compounded risk-free rate (all maturities) is 6% and the time to maturity is one year. What is the price of a one-year European put option on the stock with a strike price of $50?

$2.09

$7.52

$3.58

$9.91
Answer #1

As per put call parity
Call Price + Strike Price*e^(-rt) = Put price + Stock Price
6+50*e^(-6%*1)=Put Price+51

Put Price =6+50*e^(-6%*1)-51 =2.09

Option a is correct option